OpenStudy (anonymous):

How to solve: square root of r+4 + the square root of 2 is equal to 2?

OpenStudy (lgbasallote):

\[\sqrt{r+4} + \sqrt 2 = 2\] yes?

OpenStudy (anonymous):

put the sqrt 2 on the other side and square both sides

OpenStudy (anonymous):

square both sides. \[(\sqrt{r+4}+\sqrt{2})^{2}=2^{2}\] \[r+4+2+2(\sqrt{r+4})(\sqrt{2})=4\]...........by applying identity. \[r+6+2({r+4})({2})=4\] \[r+6+{4r+16}=4\] \[5r+20=4\] \[5r+20-20=4-20\] \[5r=-16\] \[r=\frac {-16}{5}\]

OpenStudy (anonymous):

I am working on other types of equetions, equations with two radicals. I was trying to facro 5r+16=0 to get two answers. In this case I have only one solution?? Can a solution of radical be negative?

OpenStudy (anonymous):

never

OpenStudy (anonymous):

sol. of a radical is always positive.

OpenStudy (anonymous):

so this equation has no solution?

OpenStudy (anonymous):

or if this checks into the equation i can say this satisfy the equation?